# Reconstruction of QCD first-order phase transition from neutron star measurements

### v1

@@ Abstract
The potential hadron-to-quark phase transition in neutron stars has not been fully understood as the property of cold, dense, and strongly interacting matter cannot be theoretically described by the first-principle perturbative calculations, nor have they been systematically measured through terrestrial low-to-intermediate energy heavy-ion experiments. Given the Tolman–Oppenheimer–Volkoff (TOV) equations, the equation of state (EoS) of the neutron star (NS) matter can be constrained by the observations of NS mass, radius, and tidal deformability. However, large observational uncertainties and the limited number of observations currently make it challenging to strictly reconstruct the EoS, especially to identify interesting features such as a strong first-order phase transition. In this work, we study the dependency of reconstruction quality of the phase transition on the number of NS observations of mass and radius as well as their uncertainty, based on a fiducial EoS. We conquer this challenging problem by constructing a neural network, which allows one to parameterize the EoS with minimum model-dependency, and by devising an algorithm of parameter optimization based on the analytical linear response analysis of the TOV equations. This work may pave the way for the understanding of the phase transition features in NSs using future $X$-ray and gravitational wave measurements.
@@ I Introduction
To quantitatively address this question, a Bayesian analysis must be conducted in a non-parametric manner. Specifically, the EoS should be represented by a flexible enough functional form capable of capturing the full complexity of the system, e.g., those parametrized with sufficient number of parameters. This approach guarantees that, in the ideal scenario with sufficient amount of high-precision NS observations, the EoS can be reconstructed with high fidelity, capturing intricate details such as a first-order PT. While Gaussian Processes (GPs) are another prominent example of non-parametric methods that provide flexible, probabilistic models, we opt for a deep neural network (DNN) representation of the EoS, given their scalability and ability to handle complex, high-dimensional parameter spaces efficiently, which is crucial for the large-scale nature of our analysis. A non-parametric Bayesian analysis is then only feasible if one can update the involved large set of parameters with high efficiency.
@@
Gradient-based optimization has proven successful in parameter-intensive tasks, so the gradients of NS observables with respect to EoS parameters can be leveraged for this purpose. Such a numerical tool has been partially developed in previous works by some of the authors [41, 42], where the EoS is represented by a DNN, and the mapping between the EoS and the $M$-$R$ relation (i.e., the TOV equations) is approximated by another DNN. This setup enables numerical computation of EoS parameter gradients via auto-differentiation. However, the TOV-equation-approximator DNN is not guaranteed to accurately reproduce the EoS-to-($M$-$R$) mapping, and quantifying the associated systematic uncertainties remains non-trivial. In this work, we resolve this issue by avoiding the use of the TOV-equation-approximator DNN. Instead, we solve the TOV equations directly and compute the derivatives with respect to changes in the EoS in a numerically efficient manner. This method, also referred to as physics-driven learning, has been applied to the study of other inverse problems in nuclear physics [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. For recent reviews, see [55, 56].
@@
With newly developed tool introduced in Sec. II, we systematically study the degradation of the reconstruction quality of a fiducial EoS in Sec. III, by gradually increasing the uncertainty of data points on its corresponding $M$-$R$ curve. The “lower limit” of the $M$-$R$ observations needed to verify the PT of this fiducial EoS will also be discussed. Following the summary and discussion in Sec. IV, we provide appendices for a numerical validation of the derivative analysis (A) and for details of importance sampling (B), respectively.
II Description of Algorithm
@@
The key to reconstructing the EoS from NS observables in a non-parametric manner – especially considering discrete and noisy observational data – is to perform Bayesian Inference with an arbitrary large number of parameters. This can be achieved if one can efficiently compute the derivative of the posterior distribution with respect to the model parameters. Noting that the posterior distribution takes NS observables as input, which depends on the EoS implicitly, computing the parameter gradients is not straightforward. This section provides a detailed description of our algorithm that computes the parameter gradients based on a linear response analysis of the TOV equations. The gradient is useful not only in guiding the direction of parameter update iterations, but also in physics analysis of the sensitivity of NS observables with respect to changes in the EoS.
@@ II.1 Likelihood distribution of the Equation of State
Performing BA in a high-dimensional parameter space (with $N_{\mathrm{prm}}\sim 10^{2-3}$), even if only searching for the MAP, is generally challenging. In statistics, Markov-Chain Monte-Carlo (MCMC) [66] is usually invoked to approach the region around ${\bm{\vartheta}}^{\mathrm{opt}}$ and construct the Posterior distribution. However, the MCMC procedure is only applicable to problems with a few parameters; for those with $N_{\mathrm{prm}}\gtrsim 100$, MCMC usually fails to converge within reasonable computation time.
Nevertheless, if the loss function is differentiable with respect to the model parameters, one may obtain ${\bm{\vartheta}}^{\mathrm{opt}}$ according to a gradient-based method which is applicable to problems with huge amount of parameters. Fortunately, our problem of interest falls into such a category. This is not obvious because $\chi^{2}$ is the function of the NSs’ masses, radii, and tidal deformabilities, all of which depend on the EoS implicitly. While one cannot compute the derivative of implicit functions in general, we perform a linear response analysis to the TOV equations and obtain the formulae of changes in mass and radius with respect to an arbitrary perturbation in the EoS, or in the central pressure of a NS. Thus, the parameter gradient of $\chi^{2}$ can be computed.
While details will be shown in Sec. II.2, here we outline the procedures of finding $\bm{\vartheta}_{\mathrm{opt}}$ according to the gradient-based iteration method:
  • (a)
  • (b)
    At the $n^{\mathrm{th}}$ iteration with parameter set $\bm{\vartheta}^{(n)}$, compute the EoS, solve the TOV equations and the differential equations for the variations (2628,34), and compute $\nabla_{\bm{\vartheta}}\chi^{2}|_{\bm{\vartheta}=\bm{\vartheta}^{(n)}}$ according to (51).
  • (c)
  • (d)
  • (a)
  • (b)
  • (c)
  • (d)
By repeating (a)-(d) and collecting sufficient numbers of samples, we obtain the posterior distribution of EoS. It should be noted that the true $\mathrm{P}(\bm{\vartheta})$ is not necessarily Gaussian, and one may still apply this framework to extract the EoS distribution from realistic observations of NS masses and radii, which follow non-Gaussian distribution. It is worth noting that we parameterize the EoS by always constraining the speed of sound to be positive and causal, so that any EoS sampled from the final posterior distribution is physical. Details of the EoS parameterization will be given in Sec. II.3, and we also supplement in Appendix B for a proof of the importance sampling method and its uncertainty estimation.
@@ II.2.1 Preparation: change of variables in the TOV equations
Given the central pressure $P(r=0)=P_{c}$ as well as the initial condition for mass $m(r=0)=0$, one can solve the differential equations for $P(r)$ and $m(r)$ with increasing $r$ starting from the center and progressing outward, until the pressure decreases to zero, $P(r)=0$. The final-state radius, labeled as $R$, defines the radius of a NS, and the enclosed mass is the total mass of the NS, $M=m(R)$. In practice, the boundary (surface) pressure is usually taken as a non-zero but small enough value, and we denote such a value as $P_{\text{bnd}}$.
In addition to the NS mass and radius, tidal deformability is another macroscopic observable that is sensitive to the nuclear EoS [74, 75, 76, 77]. It characterizes the response relationship of the induced mass-quadrupole moment of a NS to an external perturbing tidal field (the gravitational field of a companion star), which can be inferred from gravitational wave detections of binary NS merger events. Although our examples of the EoS reconstruction in this work only involve information of mass and radius, we show the differential equation for computing tidal deformability and sensitivity analysis for the convenience of future study.
@@ II.2.3 Perturbation in the central pressure
with $\mathcal{K}_{m,y}=\mathcal{K}_{v,y}=\mathcal{K}_{m,\kappa_{s}}=\mathcal{K}_{v,\kappa_{s}}=0$.
@@ II.2.4 Variation analysis: summary
Here we summarize the variation analysis procedures. Given perturbations in the EoS function and in the central pressure, we solve Eqs. (2628, 31) and (34), respectively, and obtain the derivatives at the $P=P_{\mathrm{bnd}}$ edge,
@@
where $\bm{\vartheta}$ is the shorthand of all the EoS parameters and the central pressures of different NSs. A numerical validation of Eqs. (2628, 31) and (34) is provided in Appendix A. The response of $\Lambda$ with respect to a perturbation in $\vartheta_{k}$ can be given by
@@
With the variations computed in an efficient manner, we are able to update the model parameters in the numerical reconstruction of the EoS. Moreover, the variational analysis is valuable by itself — it directly gives the sensitivity of the NS mass, radius, and tidal deformability with respect to the change in the equation of state. In Fig. 1, we start from the SFHo [79] equation of state, $\varepsilon=\varepsilon_{\mathrm{SFHo}}(P)$, and respectively introduce two types of perturbation in the energy density. First, a local change which can be formulated by a Dirac-$\delta$ function as in Eq. (29); second, a perturbation in the transition pressure and latent heat of a first-order phase transition at $P=P_{\mathrm{pert}}$,
@@
For a NS with central pressure $P_{c}$, we compute the change of $M$, $R$, and $\Lambda$, and scale the derivatives to show their relative deviations.
@@ II.3 Parametrization of the Equation of State
We parameterize the inverse speed of sound squared containing both regular parts and Dirac-$\delta$ functions corresponding to possible first-order phase transitions,
@@
where $\Theta$ is the step function, $\bm{\theta}$ are parameters of the regular sector, and $\xi_{l}$ indicates the log-pressure of the first-order phase transition point, and $\Delta_{l}\geq 0$ the latent heat. Keeping in mind that there could be rich phase structures, we have introduced a parametrization scheme that allows alternative number of PT points, and the number of $l$’s is a hyperparameter of the parametrization. $\mathcal{F}$ is represented as piece-wise functions accounting for the fact that the speed of sound value could be discontinuous at the phase transition point. Each $\mathcal{F}_{l}$ is parameterized as $\mathcal{F}_{l}(\xi|\bm{\theta}_{l})=1+e^{{f}_{l}(\xi|\bm{\theta}_{l})}$ to fulfill the causality condition.
@@
While what have been discussed above (including general discussions of the algorithm in Sec. II.1 and II.2) are applicable for arbitrary parametrization of the EoS, we exploit the deep neural network (DNN) as an unbiased and flexible parametrization [80, 81] of the EoS.
DNN is a parametrization scheme which can approximately express any $\mathbb{R}_{n}\to\mathbb{R}_{m}$ function mapping between independent variables (inputs) $\bm{x}=\{x_{1},\cdots,x_{n}\}$ and dependent variables (outputs) $\bm{y}=\{y_{1},\cdots,y_{m}\}$, $\bm{y}=\bm{y}(\bm{x})$. It constructs the functional form by iteratively composing $N$ simple building blocks (also called layer representing a vector-to-vector function). Each layer performs a linear transformation on the output from the preceding layer, followed by an element-wise non-linear transformation dictated by the activation function $\sigma^{(\ell)}(z)$
@@
Along these lines, DNN delivers a powerful tool to learn about essential physical quantities in nuclear physics by capturing complex, non-linear patterns that traditional models struggle with [56]. For instance, DNN helps phase transition identification in heavy-ion collisions [83, 84, 85] as well as in lattice study [86, 87], improves parton distribution function (PDF) [43, 44, 45] and spectral function reconstruction [88, 47, 48], quasi-particle modeling [49, 50], extraction of heavy quark potentials [46] and hadron-hadron interactions [52, 53] from lattice QCD data, etc. It could also help nuclear many-body calculations as an unbiased but flexible Ansatz [89, 90, 91]. In addition, DNN-based emitting source functions offer new insights into particle production in heavy-ion collisions [54].

### v2

# Towards constraining QCD phase transitions in neutron star interiors: Bayesian Inference with TOV linear response analysis

@@ Abstract
The potential hadron-to-quark phase transition in neutron stars has not been fully understood as the property of cold, dense, and strongly interacting matter cannot be theoretically described by the first-principle perturbative calculations, nor have they been systematically measured through terrestrial low-to-intermediate energy heavy-ion experiments. Given the Tolman–Oppenheimer–Volkoff (TOV) equations, the equation of state (EoS) of the neutron star (NS) matter can be constrained by the observations of NS mass, radius, and tidal deformability. However, large observational uncertainties and the limited number of observations currently make it challenging to strictly reconstruct the EoS, especially to identify interesting features such as a strong first-order phase transition. In this work, we study the dependency of reconstruction quality of the phase transition on the number of NS observations of mass and radius as well as their uncertainty, based on a fiducial EoS. We conquer this challenging problem by constructing a neural network, which allows one to parametrize the EoS with minimum model-dependency, and by devising an algorithm of parameter optimization based on the analytical linear response analysis of the TOV equations. This work may pave the way for the understanding of the phase transition features in NSs using future $x$-ray and gravitational wave measurements.
@@ I Introduction
To address this question quantitatively, one must infer the EoS based on a sufficiently flexible assumption (i.e., to perform a nonparametric Bayesian analysis) while ensuring that the update process remains efficient for handling numerous observations. Specifically, the EoS should be represented by a flexible enough functional form capable of capturing the full complexity of the system, e.g., those parametrized with sufficient number of parameters. This approach guarantees that, in the ideal scenario with sufficient amount of high-precision NS observations, the EoS can be reconstructed with high fidelity, capturing intricate details such as a first-order PT. While Gaussian Processes (GPs) are another prominent example of nonparametric methods that provide flexible, probabilistic models, we opt for a deep neural network (DNN) representation of the EoS, given their scalability and ability to handle complex, high-dimensional parameter spaces efficiently, which is crucial for the large-scale nature of our analysis. A nonparametric Bayesian analysis is then only feasible if one can update the involved large set of parameters with high efficiency.
@@
Gradient-based optimization has proven successful in parameter-intensive tasks, so the gradients of NS observables with respect to EoS parameters can be leveraged for this purpose. Such a numerical tool has been partially developed in previous works by some of the authors [41, 42], where the EoS is represented by a DNN, and the mapping between the EoS and the $M$-$R$ relation (i.e., the TOV equations) is approximated by another DNN. This setup enables numerical computation of EoS parameter gradients via auto-differentiation. However, the TOV-equation-approximator DNN is not guaranteed to accurately reproduce the EoS-to-($M$-$R$) mapping, and quantifying the associated systematic uncertainties remains nontrivial. In this work, we resolve this issue by avoiding the use of the TOV-equation-approximator DNN. Instead, we solve the TOV equations directly and compute the derivatives with respect to changes in the EoS in a numerically efficient manner. This method, also referred to as physics-driven learning, has been applied to the study of other inverse problems in nuclear physics [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. For recent reviews, see [55, 56].
@@
In this work, we develop the computational method and perform a proof-of-principle study, based on which the aforementioned question could eventually be answered. With newly developed tool introduced in Sec. II, we systematically study the degradation of the reconstruction quality of a fiducial EoS in Sec. III, by gradually increasing the uncertainty of data points on its corresponding $M$-$R$ curve. The “lower limit” of the $M$-$R$ observations needed to verify the PT of this fiducial EoS will also be discussed. Following the summary and discussion in Sec. IV, we provide appendices for a numerical validation of the derivative analysis (A) and for details of importance sampling (B), respectively.
II Description of the Algorithm
@@
The key to reconstructing the EoS from NS observables in a nonparametric manner – especially considering discrete and noisy observational data – is to perform Bayesian Inference with an arbitrary large number of parameters. This can be achieved if one can efficiently compute the derivative of the posterior distribution with respect to the model parameters. Noting that the posterior distribution takes NS observables as input, which depends on the EoS implicitly, computing the parameter gradients is not straightforward. This section provides a detailed description of our algorithm that computes the parameter gradients based on a linear response analysis of the TOV equations. The gradient is useful not only in guiding the direction of parameter update iterations, but also in physics analysis of the sensitivity of NS observables with respect to changes in the EoS.
@@ II.1 Likelihood distribution of the Equation of State
Performing BA in a high-dimensional parameter space (with $N_{\mathrm{prm}}\sim 10^{2-3}$), even if only searching for the MAP, is generally challenging. In statistics, Markov-Chain Monte-Carlo (MCMC) [66] is usually invoked to approach the region around ${\bm{\vartheta}}^{\mathrm{opt}}$ and construct the Posterior distribution. At each step of the MCMC sampling, a new parameter set is proposed randomly, and its acceptance probability is designed such that the desired likelihood function can be approached. As the acceptance rate decreases rapidly with the increasing number of parameters, the MCMC procedure is only applicable to problems with a few parameters; for those with $N_{\mathrm{prm}}\gtrsim 100$, MCMC usually fails to converge within reasonable computation time.
Nevertheless, if the loss function is differentiable with respect to the model parameters, one may obtain ${\bm{\vartheta}}^{\mathrm{opt}}$ according to gradient-based methods, which is applicable to problems with huge amount of parameters. Fortunately, our problem of interest falls into such a category. This is not obvious because $\chi^{2}$ is the function of the NSs’ masses, radii, and tidal deformabilities, all of which depend on the EoS implicitly. While one cannot compute the derivative of implicit functions in general, we perform a linear response analysis to the TOV equations and obtain the formulae of changes in mass and radius with respect to an arbitrary perturbation in the EoS, or in the central pressure of a NS. Thus, the parameter gradient of $\chi^{2}$ can be computed. With $\nabla_{\bm{\vartheta}}\chi^{2}$, one may update $\bm{\vartheta}$ iteratively to minimize $\chi^{2}$, i.e., maximize the likelihood function, and explore the region around that. By doing so, one may propose parameter sets most efficiently to perform a BA.
@@ II.2.1 Preparation: change of variables in the TOV equations
Given the central pressure $P(r=0)=P_{c}$ as well as the initial condition for mass $m(r=0)=0$, one can solve the differential equations for $P(r)$ and $m(r)$ with increasing $r$ starting from the center and progressing outward, until the pressure decreases to zero, $P(r)=0$. The final-state radius, labeled as $R$, defines the radius of a NS, and the enclosed mass is the total mass of the NS, $M=m(R)$. In practice, the boundary (surface) pressure is usually taken as a nonzero but small enough value, and we denote such a value as $P_{\text{bnd}}$.
In addition to the NS mass and radius, tidal deformability is another macroscopic observable that is sensitive to the nuclear EoS [71, 72, 73, 74]. It characterizes the response relationship of the induced mass-quadrupole moment of a NS to an external perturbing tidal field (the gravitational field of a companion star), which can be inferred from gravitational wave detections of binary NS merger events. Although our examples of the EoS reconstruction presented later in this work only involve information of mass and radius, we also perform linear response analysis for the tidal deformability for reference.
@@ II.2.2 Perturbation in the EoS
Figure 1: NS equations of state ($a$) and their corresponding $M$-$R$ ($b$) and $\Lambda$-$R$ ($c$) curves. Black curves correspond to the SFHo EoS, while red and blue curves include a step function and a Gauss function perturbation on top of that, respectively. The short colored lines connect the $M$, $R$, and $\Lambda$ values of NSs with the same central pressure before and after the perturbation.
@@
For illustrative purposes, we show in Fig. 1 how the NS mass, radius, and tidal deformability respond to the change in the EoS. On top of the baseline SFHo EoS, we respectively introduce i) a step function perturbation, and ii) a Gauss function perturbation, and then compute the induced changes in NS observables. As both types of perturbation make the EoS softer, they lead to a smaller mass, smaller radius, and greater tidal deformability for a given NS with fixed central pressure. Note that while a Gauss function perturbation only changes the $M$-$R$ and $\Lambda$-$R$ curves within a finite range of $P_{c}$, a step function perturbation changes the whole curve with $P_{c}$ above the PT pressure.
@@ II.2.3 Perturbation in the central pressure
with $\mathcal{K}_{m,y}=\mathcal{K}_{v,y}=\mathcal{K}_{m,\kappa_{s}}=\mathcal{K}_{v,\kappa_{s}}=0$. Given that different points on a $M$-$R$-$\Lambda$ curve are different in $P_{c}$, we note that the derivatives, $\frac{\partial M}{\partial P_{c}}$, $\frac{\partial R}{\partial P_{c}}$, and $\frac{\partial\Lambda}{\partial P_{c}}$, determine the slope of the curve.
@@ II.2.4 Variation analysis: summary
Here we summarize the variation analysis procedures. Given perturbations in the EoS function and in the central pressure, we solve Eqs. (2426, 29) and (32), respectively, and obtain the derivatives at the $P=P_{\mathrm{bnd}}$ edge,
@@
where $\bm{\vartheta}$ is the shorthand of all the EoS parameters and the central pressures of different NSs. A numerical validation of Eqs. (2426, 29) and (32) is provided in Appendix A. The response of $\Lambda$ with respect to a perturbation in $\vartheta_{k}$ can be given by
@@
With the variations computed in an efficient manner, we are able to update the model parameters in the numerical reconstruction of the EoS. Moreover, the variational analysis is valuable by itself — it directly gives the sensitivity of the NS mass, radius, and tidal deformability with respect to the change in the equation of state. In Fig. 2, we start from the SFHo [76] equation of state, $\varepsilon=\varepsilon_{\mathrm{SFHo}}(P)$, and respectively introduce two types of perturbation in the energy density.
First, we add a small Dirac-$\delta$ function local perturbation in the EoS, $\varepsilon(P)\to\varepsilon(P)+\delta(P-P_{\mathrm{pert}})\,\delta_{\varepsilon}/P_{\mathrm{pert}}$, and compute the change in mass for a NS with its central pressure being $P_{c}$, denoted as $\delta_{M}(P_{c})$, and the functional derivative can be obtained as $\frac{\delta M(P_{c})}{\delta\varepsilon(P)}|_{P=P_{\mathrm{pert}}}=\frac{\delta_{M}(P_{c})}{\delta_{\varepsilon}}$. Varying the position of local change ($P_{\mathrm{pert}}$) and the NS of interest ($P_{c}$), we present the dimensionless relative functional derivatives (scaled by the respective value of $M$ and $\varepsilon(P_{\mathrm{pert}})$) in Fig. 2, panel ($a$), and likewise for $R$ and $\Lambda$ in panels ($b$) and ($c$).
Second, we add a small perturbation in the transition pressure and latent heat of a first-order phase transition at $P=P_{\mathrm{pert}}$,
@@
For a NS with central pressure $P_{c}$, we compute the change of $M$, $R$, and $\Lambda$, and show their scaled, relative deviations in Fig. 2, panels ($d$-$i$).
@@ II.3 Parametrization of the Equation of State
We parametrize the inverse speed of sound squared containing both regular parts and Dirac-$\delta$ functions corresponding to possible first-order phase transitions,
@@
where $\Theta$ is the step function, $\bm{\theta}$ are parameters of the regular sector, and $\xi_{l}$ indicates the log-pressure of the first-order phase transition point, and $\Delta_{l}\geq 0$ the latent heat. Keeping in mind that there could be rich phase structures, we have introduced a parametrization scheme that allows alternative number of PT points, and the number of $l$’s is a hyperparameter of the parametrization. $\mathcal{F}$ is represented as piece-wise functions accounting for the fact that the speed of sound value could be discontinuous at the phase transition point. Each $\mathcal{F}_{l}$ is parametrized as $\mathcal{F}_{l}(\xi|\bm{\theta}_{l})=1+e^{{f}_{l}(\xi|\bm{\theta}_{l})}$ to fulfill the causality condition.
@@
Here we outline the procedures of finding $\bm{\vartheta}_{\mathrm{opt}}$ according to the gradient-based iteration method:
  • (a)
  • (b)
    At the $n^{\mathrm{th}}$ iteration with parameter set $\bm{\vartheta}^{(n)}$, compute the EoS, solve the TOV equations and the differential equations for the variations (2426, 32), and compute $\nabla_{\bm{\vartheta}}\chi^{2}|_{\bm{\vartheta}=\bm{\vartheta}^{(n)}}$ according to (49).
  • (c)
  • (d)
  • (a)
  • (b)
  • (c)
  • (d)
By repeating (a)-(d) and collecting sufficient numbers of samples, we obtain the posterior distribution of EoS. It should be noted that the true $\mathrm{P}(\bm{\vartheta})$ is not necessarily Gaussian, and one may still apply this framework to extract the EoS distribution from realistic observations of NS masses and radii, which follow nonGaussian distribution. It is worth noting that we parametrize the EoS by always constraining the speed of sound to be positive and causal, so that any EoS sampled from the final posterior distribution is physical. We supplement in Appendix B for a proof of the importance sampling method and its uncertainty estimation.
Parameterization with DNN. — While what have been discussed above (including general discussions of the algorithm in Sec. II.1 and II.2) are applicable for arbitrary parametrization of the EoS, we exploit the deep neural network (DNN) as an unbiased and flexible parametrization [80, 81] of the EoS.
DNN is a parametrization scheme which can approximately express any $\mathbb{R}_{n}\to\mathbb{R}_{m}$ function mapping between independent variables (inputs) $\bm{x}=\{x_{1},\cdots,x_{n}\}$ and dependent variables (outputs) $\bm{y}=\{y_{1},\cdots,y_{m}\}$, $\bm{y}=\bm{y}(\bm{x})$. It constructs the functional form by iteratively composing $N$ simple building blocks (also called layer representing a vector-to-vector function). Each layer performs a linear transformation on the output from the preceding layer, followed by an element-wise nonlinear transformation dictated by the activation function $\sigma^{(\ell)}(z)$
@@
Along these lines, DNN delivers a powerful tool to learn about essential physical quantities in nuclear physics by capturing complex, nonlinear patterns that traditional models struggle with [56]. For instance, DNN helps phase transition identification in heavy-ion collisions [83, 84, 85] as well as in lattice study [86, 87], improves parton distribution function (PDF) [43, 44, 45] and spectral function reconstruction [88, 47, 48], quasi-particle modeling [49, 50], extraction of heavy quark potentials [46] and hadron-hadron interactions [52, 53] from lattice QCD data, etc. It could also help nuclear many-body calculations as an unbiased but flexible Ansatz [89, 90, 91]. In addition, DNN-based emitting source functions offer new insights into particle production in heavy-ion collisions [54].